OpenStudy (anonymous):

Please factor a^2 – 5a – 20

OpenStudy (anonymous):

a=15

OpenStudy (anonymous):

Explain please?

OpenStudy (anonymous):

Trying

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Hard

OpenStudy (jhannybean):

Use the quadratic formula \[\large x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}\]\[\large a = 1 \ ,\ b=-5 \ , \ c=-20\]

OpenStudy (anonymous):

^cheats

OpenStudy (anonymous):

if you get a=15 try solving the question again

OpenStudy (jhannybean):

How am i cheating?

OpenStudy (anonymous):

I though you cant use quadratic equation

OpenStudy (jhannybean):

What's stopping you?

OpenStudy (anonymous):

Nothing just my mom told me not too

OpenStudy (anonymous):

FYI its algebra 1... ^confused me...!

OpenStudy (anonymous):

I honestly think you only solve this question with the quadratic equation

OpenStudy (anonymous):

are you sure?

OpenStudy (anonymous):

Not sure but imma trying

OpenStudy (anonymous):

no he is wrong you can i got my answer like that

OpenStudy (jhannybean):

\[\large x = \frac{-(-5) \pm \sqrt{(-5)^2 -4(1)(-20)}}{2(1)}\]\[\large \frac{5\pm \sqrt{25+80}}{2}\]\[\large \frac{5 \pm \sqrt{105}}{2}\]

OpenStudy (anonymous):

jhanney using cheats again

OpenStudy (anonymous):

So its either 7.62 or -2.62

OpenStudy (anonymous):

no pak asfand is wrong again

OpenStudy (jhannybean):

\[\large a^2 -5a -20\]\[\large (a^2-5a)-20\]complete the square. \[\large (a^2-5a+\frac{25}{4})-\frac{25}{4}-20\]\[\large (a-\frac52)^2 -\frac{105}{4}\] Now to solve for a... put -105/4 on the other side.

OpenStudy (jhannybean):

There are several ways to solve this.....